# Nonparametric bounds for vaccine effects in randomized trials

### v1

@@ Abstract
Vaccine randomized trials are typically designed to be blinded, ensuring that the estimated vaccine efficacy (VE) reflects the immunological effect of the vaccine. When blinding is broken, however, the estimated VE reflects not only the immunological effect but also behavioral effects stemming from participants’ awareness of their treatment status. Recent work has proposed alternative causal estimands to the standard VE to address this issue, but their point identification results require a strong assumption: the absence of unmeasured common causes of infection risk and participants’ belief about whether they received the vaccine. Personality traits, for example, may plausibly violate this assumption. We relax this assumption and derive nonparametric causal bounds for different types of VE. We construct these bounds using two approaches: linear programming-based and monotonicity-based methods. We further consider several possible causal structures for vaccine trials and show how the nonparametric bounds differ across these scenarios. Finally, we illustrate the performance of the proposed bounds using fully synthetic data and a semi-synthetic data example based on a COVID-19 vaccine trial.
@@ 1 Introduction
In vaccine trials, broken blinding, i.e., when participants become aware of their treatment assignment, has recently been recognized as a factor that may cause the immunological effect, typically the target of the trial, to be non-identifiable (Gabriel et al.,, 2025). Moreover, the estimated vaccine efficacy (VE) in such trials has been recently shown to be a biased estimator of the true immunological effect (Stensrud et al.,, 2024; Obolski et al.,, 2024). In the presence of broken blinding, the estimated VE reflects not only the immunological effect but also behavioral changes induced by participants’ awareness of their treatment status. For example, after COVID-19 vaccination, individuals who felt more protected were found to increase social contacts, reduce mask use, and attend larger gatherings, potentially elevating their exposure to SARS-CoV-2 (Trogen and Caplan,, 2021; Buckell et al.,, 2023).
@@
Since examples of broken blinding in vaccine trials have been documented (Lazarus et al.,, 2021), it is crucial to discuss how the true VE can be estimated in such trials. As reported in Lazarus et al., (2021), a possible reason for the blinding process to fail is that the active vaccine causes more local reaction than placebo, which could unmask the active vaccine versus placebo allocation. Another reason is that the treatment may be effective and reveal itself through noticeable improvements, allowing participants to deduce their actual assignment (Boutron et al.,, 2005).
@@
Alternative causal estimands have recently been proposed to study behavioral and immunological effects arising from broken blinding in vaccine trials. One such estimand corresponds to the VE that would be observed had participants known whether they were vaccinated or not (Stensrud et al.,, 2024; Obolski et al.,, 2024). Importantly, such an estimand is not only relevant for broken-blinding scenarios but also reflects vaccine effectiveness in real-world settings, where people are aware of their vaccination status and may adjust their behavior accordingly. A recent point-identification strategy of the suggested estimands builds on the idea of measuring trial participants belief about their vaccination status.
@@
However, these point-identification results require an arguably strong assumption: that there are no unmeasured common causes of infection risk and of belief about the treatment received. A key example of such a common cause in vaccine studies is personality traits, which may violate this assumption. For instance, optimism may lead participants to believe they received the vaccine even when they did not (“wishful thinking”, Bang,, 2016), and the same optimistic disposition may also make them more outgoing, thereby increasing their chances of exposure to infectious agents.
@@
To avoid making the assumption of no unmeasured common causes, we develop in this paper nonparametric causal bounds for different types of VE. Nonparametric bounds have been proposed in a variety of contexts. Robins, (1989) and Manski, (1990) independently derived the first informative nonparametric bounds on causal effects in studies with imperfect compliance, where the standard randomization assumption cannot be invoked. Later, Horowitz and Manski, (2000) derived nonparametric bounds for randomized experiments with missing data, relying only on observed confounders and making no assumptions about the missingness mechanism. Other contexts in which nonparametric bounds have been used include, for example, mediation analysis (Robins and Richardson,, 2010; VanderWeele,, 2010; Gabriel et al., 2023a, ) and outcome-dependent sampling with unmeasured confounding (Gabriel et al.,, 2022).
@@
We construct nonparametric bounds using two different approaches: linear programming (LP)-based and monotonicity-based methods. The LP-based approach originates in the seminal paper of Balke and Pearl, (1994), whose framework enabled bounding the average treatment effect in studies with imperfect compliance. They recognized that deriving such bounds is equivalent to solving an LP optimization problem, with the target causal parameter and the constraints implied by observed probabilities expressed as linear combinations of the same underlying parameters. In the binary treatment and binary outcome setting, Balke and Pearl further derived closed-form symbolic solutions, providing analytical expressions without the need for iterative optimization. Their results were later extended to the more general instrumental variables case (see Swanson et al.,, 2018, for a review). The original Balke–Pearl bounds can be further tightened by imposing monotonicity assumptions (Balke and Pearl,, 1994; Kuroki and Cai,, 2008), and more recent work has generalized the framework to accommodate multiple scenarios (Gabriel et al., 2023b, ; Sachs et al.,, 2023).
@@
While LP-based bounds are sharp (i.e., the tightest possible bounds under the stated assumptions), they can be wide in practice, especially when the data provide limited information about the joint distribution of unmeasured variables. As an alternative, we consider a second approach based on monotonicity assumptions, namely that both the belief variable and the outcome are monotonically related to the unmeasured variable.. To this end, we build on ideas similar to those of VanderWeele, (2008), who established sign determination results for causal effects in the presence of unmeasured confounding in observational studies.
@@
We consider several possible causal structures for a vaccine trial. Each structure consists of particular assumptions about the relationships between key trial variables, which in turn yield different nonparametric bounds for the different VEs. We begin with a baseline setting that assumes no unmeasured common causes between belief and risk of infection, thereby permitting point identification (Stensrud et al.,, 2024). Next, we relax this assumption by allowing an unmeasured common cause of belief and risk of infection. We further relax the assumptions by permitting the unmeasured common cause to also directly affect adverse events (AEs). For each structure, we first provide real-world examples, supported by existing studies and then present the corresponding LP-based and monotonicity-based bounds for the different VEs.
@@
As in illustrative example, we consider the ENSEMBLE2 trial, a randomized, double-blind, placebo-controlled study evaluating a COVID-19 vaccine in adults ($\geq 18$ years) across multiple international sites (Hardt et al.,, 2022). In a prespecified safety subset, solicited local and systemic AEs occurred in 57.3% of vaccine recipients versus 22.5% of placebo recipients, raising concerns that blinding may have been compromised. Consequently, as previously noted (Stensrud et al.,, 2024; Obolski et al.,, 2024), the estimated VE in this trial may not reflect the causal immunological effects of the vaccine. At the same time, it is not plausible to assume the absence of common causes influencing both belief and risk of infection (e.g., personality traits). Thus, the point identification approach of Stensrud et al., (2024) is not feasible here, motivating instead a partial identification strategy based on nonparametric bounds for VEs.
@@ 2.1 Notation
We begin by reviewing the framework of Stensrud et al., (2024). Let $A$ denote the randomized vaccine assignment, with $A=1$ for active vaccine and $A=0$ for placebo. In this framework, a key component is the introduction of the message $M$. Let $M\in\{-1,0,1\}$ indicate that vaccine status is blinded (M=1)(M=-1), that the individual received the message they are vaccinated $(M=1)$, or that they received the message they are unvaccinated $(M=0)$. Unlike a real-world setting, where $A=M$, blinding in a randomized trial implies that $M$ is fixed to 1-1.
A second key component is the belief $B\in\{0,1\}$, which indicates whether the individual believes they were vaccinated $(B=1)$ or not $(B=0)$. Additionally, let $S\in\{0,1\}$ be an indicator of an AE, and let $E$ represent the level of exposure to the infectious agent. The variable $E$ is introduced only to illustrate causal scenarios; we do not assume $E$ is measured, and it will not be used in the bound calculations. While the causal directed acyclic graphs (DAGs) (Pearl,, 2009) presented later on clarify the ordering of these variables in the scenarios we analyze, we note here that $A$ and $M$ are determined first, followed by $S$, then $B$, and finally the outcome $Y$.
@@ 2.2 Causal estimands
and $VE_{M}(a)=1-\frac{E(Y^{a,m=1})}{E(Y^{a,m=0})}$ for $a,m\in\{0,1\}^{2}$. These estimands correspond to the immunological effect of the vaccine under the message $m$ and the behavioral effect of the message at treatment level $a$, respectively. In practice, such four-arm trials have rarely been conducted, likely because it is ethically problematic to mislead participants about their vaccination status in ways that might influence their behavior and increase risk. For this reason, Stensrud et al., (2024) presented conditions under which data from a standard two-arm trial, augmented with measurements of $B$, can be used to identify and interpret causal contrasts as $VE(m)$ and $VE_{M}(a)$. We review the details in Section 2.3.
@@ 2.3 Point identification
Stensrud et al., (2024) presented assumptions under which $VE(m),\;m\in\{-1,0,1\}$, and $VE_{T}$ are identifiable from a two-arm trial $\mathcal{T}_{II}$ that also includes the belief $B$.
@@
These assumptions are plausible in most studies; for further discussion see Stensrud et al., (2024). The third assumption is inspired by the literature on separable effects (Robins and Richardson,, 2010; Stensrud et al.,, 2021, 2023).
@@
Under Assumptions 1-3, Proposition 1 in Stensrud et al., (2024) asserts that $E(Y^{a,m}),\;a,m\in\{0,1\}$, is identified from the two-arm trial $\mathcal{T}_{II}$ by $E(Y^{a,m})=p_{1.am}$, and therefore, $VE(m),\;m\in\{-1,0,1\}$, and $VE_{T}$ are identified by
@@
In some cases, Assumption 3 may fail but the VEs can still be point identified with additional data and alternative assumptions. Consider Figure 1(b), which presents a trial with possible broken blinding. In Figure 1(b) the occurrence of AEs $S$ plays a crucial role in the violation of Assumption 3. As demonstrated in our motivating example, the ENSEMBLE2 trial (Hardt et al.,, 2022), in this scenario AEs are more likely among the treated and those experiencing the AEs are more likely to believe they received the treatment (the path $A\rightarrow S\rightarrow B$). Moreover, the likelihood of AEs might be affected by an unmeasured variable $U$ that also affects the outcome $Y$ as illustrated by the path $S\leftarrow U\rightarrow Y$. This variable $U$ can represent, for example, general physiological frailty (e.g., immunosuppression) which may increase the risk of AEs and of infection. The path $M\rightarrow B\leftarrow S\leftarrow U\rightarrow Y$ violates Assumption 3.
@@
It can be shown (Stensrud et al.,, 2024, Proposition 2) that under Assumptions 1, 4-5, the expectations $E(Y^{a,m})$ for $a,m\in\{0,1\}^{2}$ are identified from the two-arm trial $\mathcal{T}_{II}$ even under broken blinding by $E(Y^{a,m})=\sum_{s}p_{1.ams}\gamma_{s.a}$. Therefore, $VE(m),\;m\in\{-1,0,1\}$ and $VE_{T}$ are identified by
@@ 3 Motivation for nonparametric bounds
Arguably, the strongest of the identification assumptions described above is Assumption 3. It will fail if there is an unmeasured common cause, $U$, between $B$ and $Y$, leading to the unblocked path $M\rightarrow B\leftarrow U\rightarrow Y$, as illustrated in Figure 2. These unmeasured variables $U$ can be, for example, psychological factors such as stress or illness anxiety disorder (hypochondria). In the two-arm trial, participants with higher anxiety levels (high $U$) might be more likely to believe they did not receive the vaccine ($B=0$) and are also more likely to obey the public health guidelines and minimize their exposure to the infectious agent ($E$), which can decrease the risk of infection ($Y$). Additionally, elevated anxiety levels can directly influence the immune system and susceptibility to infection through physiological mechanisms such as hormonal changes and immunosuppression (Cohen et al.,, 1991; Segerstrom and Miller,, 2004).
@@
  1. 1. Figures 1(a)-1(b).
    The path $M\rightarrow B\leftarrow U\rightarrow Y$ does not exist and thus point identification is possible.
  2. 2. Figure 2.
    The path $M\rightarrow B\leftarrow U\rightarrow Y$ exists and thus point identification is not possible. A special case of this scenario captures the scenario described in Figure 1(b), but because $S$ is unobserved, point identification is not possible.
  3. 3. Figure 3(a).
    The path $M\rightarrow B\leftarrow U\rightarrow Y$ violates Assumption 3 and thus point identification is not possible. The paths $S\rightarrow Y$ and $S\leftarrow U\rightarrow Y$ do not exist.
  4. 4. Figures 3(b), 3(c) and 3(d).
    The path $M\rightarrow B\leftarrow U\rightarrow Y$ is present, together with one or both of the following paths between $S$ and $Y$. If only $S\rightarrow Y$ exists, then Assumption 3 is violated (as in Group 3). If $S\leftarrow U\rightarrow Y$ exists, then Assumption 4 is violated. In either situation, point identification is not possible.
@@
Table 1: Summary of scenarios considered in this paper. A checkmark (✓) indicates “yes,” and a cross ($\times$) indicates “no.”
Figure Identifiable? Violating paths Path $S\rightarrow Y$
1(a)
1(b)
2 $\times$ $B\leftarrow U\rightarrow Y$
3(a) $\times$ $B\leftarrow U\rightarrow Y$ $\times$
3(b) $\times$ $B\leftarrow U\rightarrow Y$
3(c) $\times$ $B\leftarrow U\rightarrow Y$, $S\leftarrow U\rightarrow Y$ $\times$
3(d) $\times$ $B\leftarrow U\rightarrow Y$, $S\leftarrow U\rightarrow Y$
For the LP-based bounds we consider the following assumption.
@@ 2 $\times$ $B\leftarrow U\rightarrow Y$
As we show in Appendix B.1, under the nonparametric structural equation model with independent errors (NPSEM-IE) (Pearl,, 2009), Assumption 6 implies that in the two-arm trial, the following condition holds for $a,a^{\prime},m\in\{0,1\}^{3}$:
@@
In words, since $M$ affects $Y$ only through $B$, the potential outcome $Y^{a^{\prime},m}$ is identical to $Y^{a^{\prime},m=-1}$ within the subpopulation of individuals for whom $B^{a^{\prime},m=-1}=m$. Nevertheless, the implication (6) is strong: it requires the condition to hold at the individual level and thus constitutes a cross-world assumption, involving the distribution of $Y^{a^{\prime},m}$ conditional on $B^{a^{\prime},m=-1}$ for $m=0,1$. Nevertheless, as discussed later, our results remain the same even if condition (6) holds only in expectation, namely $E\!\left(Y^{a^{\prime},m}\mid A=a,B^{a^{\prime},m=-1}=m,U=u\right)=E\!\left(Y^{a^{\prime},m=-1}\mid A=a,B^{a^{\prime},m=-1}=m,U=u\right).$
The subsequent sections present LP-based and monotonicity-based bounds for Groups 2-4. In Sections 7 and 8, we compare between the LP-based or monotonicity-based bounds in terms of length, and examine the factors that affect the width of the bounds. For Group 2, under Figure 2, we explain in length the methods we leverage in this paper to construct the nonparametric, valid, and sharp bounds. Bounds are considered valid if they do not exclude any possible value of the estimand based on the true probabilities that can be estimated from the data. Bounds are sharp if every value within them represents a feasible value for the estimand, given these estimable probabilities (Jonzon et al.,, 2025).
@@
4 Nonparameric bounds for vaccine effects under Figure 2
LP bounds are based on the potential response approach; see MacLehose et al., (2005) for a gentle review of the approach. Compared to existing research, the belief $B$ introduces additional nuances in these definitions. As a result, the standard LP-based results (Balke and Pearl,, 1997) cannot be applied directly. We therefore derive LP-based bounds tailored to our setting, presenting the main ideas in the text and deferring the technical details to Appendix C.

### v2

@@ Abstract
Randomized controlled vaccine trials are typically designed to be blinded, ensuring that the estimated vaccine efficacy (VE) reflects the immunological effect of vaccination. When blinding is broken, however, the estimated VE may reflect not only the immunological effect but also behavioral effects stemming from participants’ awareness of their treatment status. Recent work has proposed alternative causal estimands to the standard VE to address this issue, but their point identification results require a strong assumption: the absence of unmeasured common causes of infection risk and participants’ belief about whether they received the vaccine. Personality traits, for example, may plausibly violate this assumption. We relax this assumption and derive nonparametric causal bounds for different types of VE. We construct these bounds using two approaches: linear programming-based and monotonicity-based methods. We further consider several possible causal structures for vaccine trials and show how the nonparametric bounds differ across these scenarios. Finally, we illustrate the performance of the proposed bounds using fully synthetic data and a semi-synthetic data example based on a COVID-19 vaccine trial.
@@ 1 Introduction
In vaccine trials, broken blinding, i.e., when participants become aware of their treatment assignment, has recently been recognized as a factor that may cause the immunological effect, typically the target of the trial, to be non-identifiable (Gabriel et al., 2025). Moreover, the estimated vaccine efficacy (VE) in such trials has been recently shown to be a biased estimator of the true immunological effect (Stensrud et al., 2024; Obolski et al., 2024). In the presence of broken blinding, the estimated VE reflects not only the immunological effect but also behavioral changes induced by participants’ awareness of their treatment status. For example, after COVID-19 vaccination, individuals who felt more protected were found to increase social contacts, reduce mask use, and attend larger gatherings, potentially elevating their exposure to SARS-CoV-2 (Trogen and Caplan, 2021; Buckell et al., 2023).
@@
Since examples of broken blinding in vaccine trials have been documented (Lazarus et al., 2021), it is crucial to discuss how the VE can be estimated. As reported in Lazarus et al. (2021), a possible reason for the blinding process to fail is that the active vaccine causes more local reaction than placebo, which could unmask the active vaccine versus placebo allocation.
@@
Alternative causal estimands have recently been proposed to study behavioral and immunological effects arising from broken blinding in vaccine trials. One such estimand corresponds to the VE that would be observed had participants known whether they were vaccinated or not (Stensrud et al., 2024; Obolski et al., 2024). Importantly, such an estimand is not only relevant for broken-blinding scenarios but also reflects vaccine effectiveness in real-world settings, where people are aware of their vaccination status and may adjust their behavior accordingly. The point-identification strategy of the suggested estimands builds on the idea of measuring trial participants belief about their vaccination status.
@@
However, these point-identification results require an arguably strong assumption: that there are no unmeasured common causes of infection risk and of belief about the treatment received. A key example of such a common cause in vaccine studies is personality traits, which may violate this assumption. For instance, optimism may lead participants to believe they received the vaccine even when they did not (“wishful thinking”, Bang, 2016), and the same optimistic disposition may also make them more outgoing, thereby increasing their chances of exposure to infectious agents.
@@
We develop in this paper nonparametric causal bounds for different types of VE, that avoid making the assumption of no unmeasured common causes. Nonparametric bounds have been proposed in a variety of contexts: Robins (1989) and Manski (1990) independently derived the first informative nonparametric bounds on causal effects in studies with imperfect compliance, where the standard randomization assumption cannot be invoked. Later, Horowitz and Manski (2000) derived nonparametric bounds for randomized experiments with missing data, relying only on observed confounders and making no assumptions about the missingness mechanism. Other contexts in which nonparametric bounds have been used include, for example, mediation analysis (Robins and Richardson, 2010; VanderWeele, 2010; Gabriel et al., 2023a) and outcome-dependent sampling with unmeasured confounding (Gabriel et al., 2022).
@@
We construct nonparametric bounds using two different approaches: linear programming (LP)-based and monotonicity-based methods. The LP-based approach originates in the seminal paper of Balke and Pearl (1994), whose framework enabled bounding the average treatment effect in studies with imperfect compliance. They recognized that deriving such bounds is equivalent to solving an LP optimization problem, with the target causal parameter and the constraints implied by observed probabilities expressed as linear combinations of the same underlying parameters. In the binary treatment and binary outcome setting, Balke and Pearl further derived closed-form symbolic solutions, providing analytical expressions without the need for iterative optimization. Their results were later extended to the more general instrumental variables case (see Swanson et al., 2018, for a review). The original Balke–Pearl bounds can be further tightened by imposing monotonicity assumptions (Balke and Pearl, 1994; Kuroki and Cai, 2008), and more recent work has generalized the framework to accommodate multiple scenarios (Gabriel et al., 2023b; Sachs et al., 2023).
@@
While LP-based bounds are sharp (i.e., the tightest possible bounds under the stated assumptions), they can be wide in practice, especially when the data provide limited information about the joint distribution of unmeasured variables. As an alternative, we consider a second approach based on monotonicity assumptions, namely that both the belief variable and the outcome are monotonically related to the unmeasured variable. To this end, we build on ideas similar to those of VanderWeele (2008), who established sign determination results for causal effects in the presence of unmeasured confounding in observational studies.
@@
We consider several possible causal structures for a vaccine trial. Each structure consists of particular assumptions about the relationships between key trial variables, which in turn yield different nonparametric bounds for the different VEs. We begin with a baseline setting that assumes no unmeasured common causes between belief and risk of infection, thereby permitting point identification (Stensrud et al., 2024). Next, we relax this assumption by allowing an unmeasured common cause of belief and risk of infection. We further relax the assumptions by permitting the unmeasured common cause to also directly affect adverse events (AEs). For each structure, we first provide real-world examples, supported by existing studies and then present the corresponding LP-based and monotonicity-based bounds for the different VEs.
@@
As in illustrative example, we consider the ENSEMBLE2 trial, a randomized, double-blind, placebo-controlled study evaluating a COVID-19 vaccine in adults ($\geq 18$ years) across multiple international sites (Hardt et al., 2022). In a prespecified safety subset, solicited local and systemic AEs occurred in 57.3% of vaccine recipients versus 22.5% of placebo recipients, raising concerns that blinding may have been compromised. Consequently, as previously noted (Stensrud et al., 2024; Obolski et al., 2024), the estimated VE in this trial may not reflect the causal immunological effects of the vaccine. At the same time, it is not plausible to assume the absence of common causes influencing both belief and risk of infection (e.g., personality traits). Thus, the point identification approach of Stensrud et al. (2024) is not feasible here, motivating instead a partial identification strategy based on nonparametric bounds for VEs.
@@ 2.1 Notation
We begin by reviewing the framework of Stensrud et al. (2024). Let $A$ denote the randomized vaccine assignment, with $A=1$ for active vaccine and $A=0$ for placebo. In this framework, a key component is the introduction of the message $M$. Let $M\in\{-1,0,1\}$ indicate that vaccine status is blinded (M=1)(M=-1), that the individual received the message they are vaccinated $(M=1)$, or that they received the message they are unvaccinated $(M=0)$. Unlike a real-world setting, where $A=M$, blinding in a randomized trial implies that $M$ is fixed to 1-1.
A second key component is the belief $B\in\{0,1\}$, which indicates whether the individual believes they were vaccinated $(B=1)$ or not $(B=0)$. Additionally, let $S\in\{0,1\}$ be an indicator of an AE, and let $E$ represent the level of exposure to the infectious agent. The variable $E$ is introduced only to illustrate causal scenarios; we do not assume $E$ is measured, and it will not be used in the bound calculations. While the causal directed acyclic graphs (DAGs) (Pearl, 2009) presented later on clarify the ordering of these variables in the scenarios we analyze, we note here that $A$ and $M$ are determined first, followed by $S$, then $B$, and finally the outcome $Y$.
@@ 2.2 Causal estimands
and $VE_{M}(a)=1-\frac{E(Y^{a,m=1})}{E(Y^{a,m=0})}$ for $a,m\in\{0,1\}^{2}$. These estimands correspond to the immunological effect of the vaccine under the message $m$ and the behavioral effect of the message at treatment level $a$, respectively. In practice, such four-arm trials have rarely been conducted, likely because it is ethically problematic to mislead participants about their vaccination status in ways that might influence their behavior and increase risk. For this reason, Stensrud et al. (2024) presented conditions under which data from a standard two-arm trial, augmented with measurements of $B$, can be used to identify and interpret causal contrasts as $VE(m)$ and $VE_{M}(a)$. We review the details in Section 2.3.
@@ 2.3 Point identification
Stensrud et al. (2024) presented assumptions under which $VE(m),\;m\in\{-1,0,1\}$, and $VE_{T}$ are identifiable from a two-arm trial $\mathcal{T}_{II}$ that also includes the belief $B$.
@@
These assumptions are plausible in most studies; for further discussion see Stensrud et al. (2024). The third assumption is inspired by the literature on separable effects (Robins and Richardson, 2010; Stensrud et al., 2021, 2023).
@@
Under Assumptions 1-3, Proposition 1 in Stensrud et al. (2024) asserts that $E(Y^{a,m}),\;a,m\in\{0,1\}$, is identified from the two-arm trial $\mathcal{T}_{II}$ by $E(Y^{a,m})=p_{1.am}$, and therefore, $VE(m),\;m\in\{-1,0,1\}$, and $VE_{T}$ are identified by
@@
In some cases, Assumption 3 may fail but the VEs can still be point identified with additional data and alternative assumptions. Consider Figure 1(b), which presents a trial with possible broken blinding. In Figure 1(b), the occurrence of AEs plays a crucial role in the violation of Assumption 3. This structure is motivated by our illustrative example, the ENSEMBLE2 trial (Hardt et al., 2022), in which AEs are more likely among the treated and experiencing an AE increases the likelihood of believing one received the treatment (the path $A\rightarrow S\rightarrow B$). Moreover, the likelihood of AEs might be affected by an unmeasured variable $U$ that also affects the outcome $Y$ as illustrated by the path $S\leftarrow U\rightarrow Y$. This variable $U$ can represent, for example, general physiological frailty (e.g., immunosuppression) which may increase the risk of AEs and of infection. The path $M\rightarrow B\leftarrow S\leftarrow U\rightarrow Y$ violates Assumption 3.
@@
It can be shown (Stensrud et al., 2024, Proposition 2) that under Assumptions 1, 4-5, the expectations $E(Y^{a,m})$ for $a,m\in\{0,1\}^{2}$ are identified from the two-arm trial $\mathcal{T}_{II}$ even under broken blinding by $E(Y^{a,m})=\sum_{s}p_{1.ams}\gamma_{s.a}$. Therefore, $VE(m),\;m\in\{-1,0,1\}$ and $VE_{T}$ are identified by
@@ 3 Motivation for nonparametric bounds
Arguably, the strongest of the identification assumptions described above is Assumption 3. It will fail if there is an unmeasured common cause, $U$, between $B$ and $Y$, leading to the unblocked path $M\rightarrow B\leftarrow U\rightarrow Y$, as illustrated in Figure 2. These unmeasured variables $U$ can be, for example, psychological factors such as stress or illness anxiety disorder (hypochondria). In the two-arm trial, participants with higher anxiety levels (high $U$) might be more likely to believe they did not receive the vaccine ($B=0$) and are also more likely to obey the public health guidelines and minimize their exposure to the infectious agent ($E$), which can decrease the risk of infection ($Y$). Additionally, elevated anxiety levels can directly influence the immune system and susceptibility to infection through physiological mechanisms such as hormonal changes and immunosuppression (Cohen et al., 1991; Segerstrom and Miller, 2004).
@@
  1. 1.
    Figures 11(a)-11(b). The path $M\rightarrow B\leftarrow U\rightarrow Y$ does not exist and thus point identification is possible.
  2. 2.
    Figure 2. The path $M\rightarrow B\leftarrow U\rightarrow Y$ exists and thus point identification is not possible. A special case of this scenario captures the scenario described in Figure 1(b), but because $S$ is unobserved, point identification is not possible.
  3. 3.
    Figure 3(a). The path $M\rightarrow B\leftarrow U\rightarrow Y$ violates Assumption 3 and thus point identification is not possible. The paths $S\rightarrow Y$ and $S\leftarrow U\rightarrow Y$ do not exist.
  4. 4.
    Figures 3(b), 3(c) and 3(d). The path $M\rightarrow B\leftarrow U\rightarrow Y$ is present, together with one or both of the following paths between $S$ and $Y$. If only $S\rightarrow Y$ exists, then Assumption 3 is violated (as in Group 3). If $S\leftarrow U\rightarrow Y$ exists, then Assumption 4 is violated. In either situation, point identification is not possible.
@@
The scenarios above are presented without additional measured baseline covariates beyond $S$. In practice, vaccine trials typically collect baseline covariates (other than $S$), denoted by $L$. In Appendix N, we examine when incorporating $L$ can lead to narrower bounds and provide the corresponding bounds under those settings.
To construct the bounds, we consider the following assumption.
@@
As we show in Appendix B.1, under the nonparametric structural equation model with independent errors (NPSEM-IE) (Pearl, 2009), Assumption 6 implies that in the two-arm trial, the following condition holds for $a,a^{\prime},m\in\{0,1\}^{3}$:
@@
In words, since $M$ affects $Y$ only through $B$, the potential outcome $Y^{a^{\prime},m}$ is identical to $Y^{a^{\prime},m=-1}$ within the subpopulation of individuals for whom $B^{a^{\prime},m=-1}=m$. Nevertheless, the implication of Equation (6) is strong: it requires the condition to hold at the individual level and thus constitutes a cross-world assumption, involving the distribution of $Y^{a^{\prime},m}$ conditional on $B^{a^{\prime},m=-1}$ for $m=0,1$. Nevertheless, as discussed later, our results remain the same even if condition (6) holds only in expectation, namely $E\!\left(Y^{a^{\prime},m}\mid A=a,B^{a^{\prime},m=-1}=m,U=u\right)=E\!\left(Y^{a^{\prime},m=-1}\mid A=a,B^{a^{\prime},m=-1}=m,U=u\right).$
The subsequent sections present LP-based and monotonicity-based bounds for Groups 2–4. All proofs are provided in the Appendix. In Sections 7 and 8, we compare the LP-based and monotonicity-based bounds in terms of length, and examine the factors that affect the width of the bounds. For Group 2, we explain in length the methods we leverage in this paper to construct the nonparametric bounds.
@@
4 Nonparametric bounds for vaccine effects under Figure 2
LP bounds are based on the potential response approach; see MacLehose et al. (2005) for a review of the approach. Compared to existing research, the belief $B$ introduces additional nuances in these definitions. As a result, the standard LP-based results (Balke and Pearl, 1997) cannot be applied directly. We therefore derive LP-based bounds tailored to our setting, presenting the main ideas in the text and deferring the technical details to Appendix C.